OpenStudy (experimentx):

Evaluate: given that b > a \[ \int_0^\infty {x^{a-1} \over 1 + x^b}\; dx\]

OpenStudy (anonymous):

*

OpenStudy (experimentx):

man ... I stuck at finding the simple residue.

OpenStudy (anonymous):

if my memory serves me, it is the numerator divided by the derivative of the denominator evaluated at the pole

OpenStudy (experimentx):

yeah you are correct ... had trouble fixing. \[ \huge \lim_{z \rightarrow e^{i \pi \over b}} {(z - e^{i{\pi }\over b}) z^{a-1} \over 1 + z^b} = \lim_{z \rightarrow e^{i \pi \over b}} {z^{a-1} \over -bz^{b-1}}\]

OpenStudy (experimentx):

 somehow it got this fixed. I had this Q on my exam paper ... Man i couldn't do it.